OpenStudy (sarahc):

Simplify 4 x 5 – 4 2 x 6 8 16 36 48

OpenStudy (sarahc):

**4 / 2

OpenStudy (ghazi):

4*5-4/2*6 20-4*3 20-12 8

OpenStudy (sarahc):

Thank you!!

OpenStudy (ghazi):

\[4*5-\frac{ 4 }{ 2 }*6=20-4*3=20-12=8\]

OpenStudy (ghazi):

:)

OpenStudy (sarahc):

ohh okay now i know what i did wrong. Your awesome(:

OpenStudy (ghazi):

thanks and you are welcome :D